OpenStudy (anonymous):

7(x -2) + 3(x + 2) = 5(x - 6)

OpenStudy (fibonaccichick666):

do you know how to distribute?

OpenStudy (anonymous):

yeah but this question is really hard and can not figure it out because when I did i must have done it wrong

OpenStudy (fibonaccichick666):

ok, do it step by step for me? I'll fix any errors, just describe what you are doing in each step

OpenStudy (anonymous):

7(x-2)+3(x+2)=5(x-6) 7(x + -2) + 3(x + 2) = 5(x + -6) -14 + 7x + 3(2 + x) = 5(x + -6) -14 + 6 + 7x + 3x = 5(x + -6) -14 + 6 = -8 + 7x + 3x = 5(x + -6) -8 + 10x = -30 + 5x -8 + 5x = -30 -30 + 8 = -22 5x = -22 Divide each side by '5'. x = -4.4

OpenStudy (fibonaccichick666):

ok, don't rearrange in the same step, it makes it much more confusing on you

OpenStudy (anonymous):

is it right tho and the program I use is telling him different or did I do something wrong lol

OpenStudy (fibonaccichick666):

but it looks right

OpenStudy (anonymous):

oh okay I will message my teacher and ask him

OpenStudy (fibonaccichick666):

but, -4.4 seems an odd answer

OpenStudy (anonymous):

y it is telling me its wrong

OpenStudy (fibonaccichick666):

did you round?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

-22 divided by 5 is -4.4 exactly

OpenStudy (fibonaccichick666):

wait, I see here: -8 + 10x = -30 + 5x -8 + 5x = -30 -30 + 8 = -22 5x = -22

OpenStudy (anonymous):

um idk?

OpenStudy (fibonaccichick666):

hmm wait no it's not what I thought there, I am going to do it myself, I must be missing your error give me a sec

OpenStudy (anonymous):

okay

OpenStudy (fibonaccichick666):

\[7(x -2) + 3(x + 2) = 5(x - 6)\]\[7x-14+3x+6=5x-30\]\[10x-8=5x-30\]\[5x-8=-30\]\[5x=-22\]\[x=\frac{-22}{5}\] maybe try inputting the fraction?

OpenStudy (anonymous):

i will try that too ty

OpenStudy (fibonaccichick666):

np, sorry, your hw system sucks

OpenStudy (fibonaccichick666):

@satellite73 , can you check me and him please? We are getting the same answer, but it is wrong?